# Towards Multi-Modal Animal Pose Estimation: An In-Depth Analysis

### v1

@@
Abstract.
Animal pose estimation (APE) aims to locate the animal body parts using a diverse array of sensor and modality inputs, which is crucial for research across neuroscience, biomechanics, and veterinary medicine. By evaluating 178 papers since 2013, APE methods are categorised by sensor and modality types, learning paradigms, experimental setup, and application domains, presenting detailed analyses of current trends, challenges, and future directions in single- and multi-modality APE systems. The analysis also highlights the transition between human and animal pose estimation. Additionally, 2D and 3D APE datasets and evaluation metrics based on different sensors and modalities are provided. A regularly updated project page is provided in this GitHub link.
Survey of animal pose estimation, human pose estimation, 2D and 3D pose estimation, unimodal and multi-modal pose estimation, deep learning-based pose estimation, pose estimation datasets, pose estimation metrics
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; 2024; price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: General and reference Surveys and overviews
1. Introduction
Animal pose estimation (APE) identifies the spatial configuration of an animal’s body parts using a diverse array of sensor and modality inputs, which is important in numerous disciplines (Karashchuk et al., 2021). For instance, neuroscientists rely on accurate measurements of animal movement to correlate with brain dynamics (Mathis et al., 2018; Krakauer et al., 2017; Pereira et al., 2020; Luxem et al., 2022). Biomechanists analyse the movement of specific animal body structures to comprehend biomechanical phenomena, such as mechanical properties, stress-strain relationships, and functionality of different body structures (Patel and Braae, 2013; Joska et al., 2021; Patel and Braae, 2014). In addition, ecologists monitor wildlife behaviours to develop more effective conservation strategies. Similarly, ethologists examine detailed animal gestures and postures to study intra- and inter-species communication and societal structures (Zuffi et al., 2019a; Ng et al., 2022a). Moreover, veterinary practitioners employ precise APE techniques to diagnose health conditions (Farahnakian et al., 2021; Hu et al., 2023) and assess recovery post-treatment, thereby mitigating potential disorders (Fang et al., 2021).
@@ 1. Introduction
1.1. Human Pose Estimation
@@
The field of HPE has rapidly evolved from manual feature engineering to automated feature learning over the past decade, due to the developments in deep learning approaches. Modern HPE methods are mainly RGB-based and predominantly employ convolutional neural networks (CNNs). In the early stages, models like VGG (Simonyan and Zisserman, 2014), ResNet-50 (He et al., 2016), ResNet-101 (He et al., 2016), and ResNet-152 (He et al., 2016) were employed as backbones for HPE from images or videos. The stacked hourglass network (Newell et al., 2016) then presented a novel architecture characterized by repeated bottom-up and top-down processing. This design aimed to capture and consolidate features at multiple resolutions, enabling precise localization of keypoints across scales. Subsequent innovations, such as the High-Resolution Network (HRNet) (Wang et al., 2020) and its variants (Cheng et al., 2020; Yu et al., 2021a), have further advanced spatial information preservation by maintaining high-resolution representations throughout the entire network. Recently, transformer-based methods (Yang et al., 2021; Yuan et al., 2021) have integrated attention mechanisms to capture long-range dependencies and spatial correlations among joints and appearances, expanding the possibilities of HPE. Beyond vision-based HPE, alternative sensors like inertial measurement units (IMUs) (Huang et al., 2020), millimeter-wave (mmWave) radar (Sengupta et al., 2020; Xue et al., 2021; Xie et al., 2023) and thermal cameras (Smith et al., 2023; Chen et al., 2020) have been explored, tailoring HPE to diverse use cases. Additionally, multi-sensor and multi-modal datasets (Ma et al., 2024; Yan et al., 2023; An et al., 2022) and approaches (Fürst et al., 2021; Zheng et al., 2022; An et al., 2022) have emerged as an innovative step over unimodal HPE to overcome individual sensor and modality limitations, offering a synergistic fusion of data that promises increased robustness and precision, especially in scenarios marked by occlusions, changing lighting conditions, and dynamic environments.
1.2. Definition of Animal Pose Estimation
@@
APE, as discussed in this review, is defined as the task of identifying and localizing body joints of terrestrial mammals using a diverse array of sensor inputs. This task results in the reconstruction of the animal’s pose, depicted as either a skeletal representation marked by keypoints (LABEL:fig:Skeleton_Pose_Representations) or as a comprehensive body mesh (LABEL:fig:Body_Mesh_Reconstructions).
@@
Figure 2. Transfering human pose estimation knowledge to proximal animal classes (e.g. Chimpanzees) for their pose estimation (Sanakoyeu et al., 2020).
[…]
Chiampanzee and Human
@@
Exploring APE through various sensory and modality inputs is important as animals may reside in diverse habitats which require different sensors for data collection. Most current APE methodologies are RGB-based, which predict the articulated joint positions from images or videos. However, RGB-based methods fail in many scenarios, such as nocturnal monitoring. For example, rodents are more active in the dark, and standard RGB cameras become inadequate due to poor lighting conditions. Consequently, infrared sensors have emerged as suitable tools for capturing the nocturnal activities of animals (Wang et al., 2018; Patel et al., 2023). Additionally, mmWave radar can detect movements but suffers from the scarcity of its generated point clouds; LiDAR adds value by providing depth information. Inertial sensors (Patel et al., 2017) can also track movement and orientation, offering invaluable insights but requiring on-animal attachment. In specific scenarios, even auditory cues (Li et al., 2022) can indicate animal movement. This paper also explores the potential of using multiple sensors (i.e. RGB + acoustic) to improve APE performance.
@@
There are two representations for the final estimated animal poses regardless of our sensors or modalities of the APE input data. The first encompasses either two-dimensional (2D) or three-dimensional (3D) keypoint pose representations (Karashchuk et al., 2021; Li and Lee, 2023; Nath et al., 2019; Desai et al., 2023) (LABEL:fig:Skeleton_Pose_Representations). In this category, the animal’s anatomy is represented as a tree of interconnected keypoints. These keypoints present animal body joints, and their interconnections reflect the natural relationship between these joints, providing a skeletal representation of the animal’s posture. The second method, body mesh reconstruction (Zuffi et al., 2017; Rüegg et al., 2022; Wu et al., 2023b, a; Yang et al., 2022a), presents a 3D view of the animal’s morphology (LABEL:fig:Body_Mesh_Reconstructions). Within this framework, we differentiate between model-based and model-free reconstructions. The model-based technique (e.g. SMAL (Zuffi et al., 2017)) utilizes supervision from the predefined template shape of the creature learned from 3D toy animal scans. This template is then refined by optimizing the shape and pose parameters. In particular, the shape parameters capture variations in body metrics (e.g. proportions, height, and weight), while the pose parameters address the observable deformations due to movement. The model-based approach can generate detailed models, but collecting the requisite data for new types of animals can be time-intensive, restricting their applicability to rare species. In contrast, the model-free approach operates without any supervision from the predefined template shape of the creature (e.g. MagicPony (Wu et al., 2023b), DOVE (Wu et al., 2023a) and 3D-Fauna (Li et al., 2024b)), which are more capable when estimating the animal poses for the diverse animal species.
@@
The transition from HPE to APE builds on HPE frameworks. Due to skeletal similarities between humans and certain animals like quadrupeds, HPE models for skeletal keypoint detection (Wang et al., 2020; Insafutdinov et al., 2016; Deb et al., 2023) and body mesh reconstruction (Loper et al., 2023) can be pertinent to APE, with modifications to cater to distinct anatomical and dynamic variations across animal species (Li and Lee, 2023; Rashid et al., 2017; Zuffi et al., 2017; Rüegg et al., 2022). One notable example is (Sanakoyeu et al., 2020), which involves transferring human DensePose knowledge to proximal animal classes (fig. 2). This approach leverages the DensePose-COCO dataset, which contains dense human body annotations, and adapts it to animal species like chimpanzees. The transfer is facilitated by establishing correspondences between the 3D models of humans and animals using semantic alignment, so that the learned dense pose mappings for humans can be applied to the animal class. This enables pose extraction without dense animal-specific annotations. In addition, innovations in APE that allow adaptation to unseen animals and environmental changes can also improve the robustness of HPE, particularly in challenging scenarios like emergency response, where the surrounding environment changes rapidly. However, bridging these two fields requires adaptation of knowledge and consideration of the distinct challenges and ethical issues in human and animal studies, which we will discuss next.
1.3. Challenges in Human Pose Estimation and Animal Pose Estimation
@@
There are also unique challenges in APE. The animal kingdom has a much greater morphological diversity and inter/intra-species variations than humans. Different animal species exhibit significant domain discrepancies in physical characteristics, body structures, and postures. Even within the same species, there can be a large variety in body size, shape, texture, motion, and posture, presenting difficulties in creating universally applicable models (Rüegg et al., 2022). For example, the pelage of porcupines can vary a lot, and this is often more pronounced and thicker in animals than in humans’ hair, which can make tracking skeleton segments tricky. Furthermore, many animals exhibit natural camouflage, travel in herds, interact frequently, or share similar appearances within their species (Zuffi et al., 2019a). These characteristics make it hard to distinguish individual animals and accurately estimate their poses as they blend into the background or appear similar to one another.
Environmental factors also play a pronounced role. Vegetation or terrain features in the natural environments may obstruct the view of animals. One example can be found in the AcinoSet dataset (Joska et al., 2021), where the cheetahs’ paws were excluded as the grass occluded them in most of the videos, making it difficult to estimate animal poses accurately. Moreover, changing lighting conditions can also heavily affect an animal’s appearance, especially when the sun’s position is moving in outdoor environments.
@@
Moreover, there are very limited annotated APE datasets (Yu et al., 2021b). Unlike HPE, which predominantly leans on supervised (Wang et al., 2020; Cheng et al., 2020) and semi-supervised learning methodologies (Ukita and Uematsu, 2018; Pavllo et al., 2019; Mitra et al., 2020) sustained by large-scale human pose datasets such as MSCOCO (Lin et al., 2014a) and MPII (Andriluka et al., 2014), APE is hindered by a lack of such datasets, especially for endangered species and wildlife (Zuffi et al., 2019a). Also, it is often impractical to place animals in controlled environments to obtain ground truth poses, although zoos provide a semi-structured environment towards data collection. Therefore, unsupervised approaches that do not rely on labelled data have become popular in APE (Wu et al., 2023b; Li and Lee, 2021b).
1.4. Applications of Animal Pose Estimation
@@
1.4.1. Tracking
APE has been applied in various tracking contexts; for example, in the CATER (Haalck et al., 2023) framework, APE is combined with environment reconstruction to track ants in complex natural environments, providing fine-scale motion trajectories even in visually challenging conditions. The SMART-BARN (Nagy et al., 2023) system uses APE to track multiple animal species simultaneously in a large 3D arena and achieve accurate tracking and behavioural analysis in indoor and naturalistic environments. In the FaceMap (Syeda et al., 2024) framework, APE is used for tracking the orofacial movements of mice, enabling detailed analysis of their facial expressions and related neural activities. For honey bees, APE tracked an entire colony over extended periods, capturing complex social behaviours and interactions without needing physical tags, enhancing our understanding of colony dynamics and individual behaviours (Bozek et al., 2021).
Table 1. Table listing various applications of animal pose estimation. Behav. denotes Behaviour Analysis.
@@ 1.4.1. Tracking
1 2022 I-MuPPET (Urs Waldmann and Goldlücke, 2022) Pigeon Tracking
2 2023 Bozek et al. (Bozek et al., 2021) Honey Bee Tracking
3 2023 CATER (Haalck et al., 2023) Desert Ants Tracking
4 2023 SMART-BARN (Nagy et al., 2023) Insects, Birds, Tracking
Mammals
5 2023 3D-POP (Naik et al., 2023) Birds Tracking
6 2023 Facemap (Syeda et al., 2024) Mice Tracking
7 2023 DeepWild (Wiltshire et al., 2023) Chimpanzees, Tracking
@@
8 2024 3D-MuPPET (Urs Waldmann and Kano, 2024) Pigeon Tracking
9 2023 Stathopoulos et al. (Stathopoulos et al., 2023) Sheep, Cows, Bear Articulation
@@
10 2022 Sheppard et al. (Sheppard et al., 2022) Mouse Behav.
11 2023 Voloh et al. (Voloh et al., 2023) Macaque Behav.
12 2023 BARN (Yang et al., 2023a) Macaque Behav.
13 2023 Koger et al. (Koger et al., 2023) Gelada Monkeys Behav.
14 2023 AmadeusGPT (Ye et al., 2023) Mammals Behav.
15 2024 Testard et al. (Testard et al., 2024) Macaque Behav.
1.4.2. Articulation
The authors of (Stathopoulos et al., 2023) explore how APE can be used to understand animal articulation by predicting their 3D shapes from monocular images. They demonstrate that training a 2D keypoint estimator on a small set of labelled images (50-150) can generate pseudo-labels on a more extensive set of unlabeled web images. Then, these pseudo-labels are used to train models for 3D shape reconstruction of various articulated quadrupeds.
1.4.3. Behaviour Analysis
In agriculture, animal behaviour analysis based on APE plays an important role by enabling effective livestock management and health surveillance, which in turn provides continuous monitoring of movement and also early detection of animal’s illness, thereby ensuring optimal animal welfare (Fang et al., 2021; Russello et al., 2022; Farahnakian et al., 2021; Gong et al., 2022; Doornweerd et al., 2021; Fang et al., 2022). Recently, more APE methods have been applied to study the behaviour of more species, such as macaques (Voloh et al., 2023; Yang et al., 2023a; Testard et al., 2024), Gelada monkeys (Koger et al., 2023), and mice (Sheppard et al., 2022). Since 2023, there has been a notable shift towards integrating language prompts with APE in behaviour analysis research, exemplified by developing models like AmadeusGPT (Ye et al., 2023). This combination has opened up new possibilities for more comprehensive interpretations of animal behaviour.
1.5. Scope of this Survey

### v2

@@
Abstract
Animal pose estimation (APE) aims to locate the animal body parts using a diverse array of sensor and modality inputs (e.g. RGB cameras, LiDAR, infrared, IMU, acoustic and language cues), which is crucial for research across neuroscience, biomechanics, and veterinary medicine. By evaluating 176 papers since 2011, APE methods are categorised by their input sensor and modality types, output forms, learning paradigms, experimental setup, and application domains, presenting detailed analyses of current trends, challenges, and future directions in single- and multi-modality APE systems. The analysis also highlights the transition between human and animal pose estimation, and how innovations in APE can reciprocally enrich human pose estimation and the broader machine learning paradigm. Additionally, 2D and 3D APE datasets and evaluation metrics based on different sensors and modalities are provided. A regularly updated project page is provided in this GitHub link.
keywords:
Animal pose estimation, human pose estimation, 2D and 3D pose estimation, unimodal and multi-modal pose estimation, multi-modal learning, sensor fusion, deep learning, pose estimation datasets and metrics, literature survey.
1 Introduction
Animal pose estimation (APE) identifies the spatial configuration of an animal’s body parts using a diverse array of sensor and modality inputs, which is important in numerous disciplines [55]. For instance, neuroscientists rely on accurate measurements of animal movement to correlate with brain dynamics [90, 62, 113, 85]. Biomechanists analyse the movement of specific animal body structures to comprehend biomechanical phenomena, such as mechanical properties, stress-strain relationships, and functionality of different body structures [106, 53, 107]. In addition, ecologists monitor wildlife behaviours to develop more effective conservation strategies. Similarly, ethologists examine detailed animal gestures and postures to study intra- and inter-species communication and societal structures [176, 103]. Moreover, veterinary practitioners employ precise APE techniques to diagnose health conditions [34, 47] and assess recovery post-treatment, thereby mitigating potential disorders [32].
@@ 1 Introduction
1.1 Human Pose Estimation
@@
The field of HPE has rapidly evolved from manual feature engineering to automated feature learning over the past decade, due to the developments in deep learning approaches. Modern HPE methods are mainly RGB-based and predominantly employ convolutional neural networks (CNNs). In the early stages, models like VGG [129], ResNet-50 [46], ResNet-101 [46], and ResNet-152 [46] were employed as backbones for HPE from images or videos. The stacked hourglass network [102] then presented a novel architecture characterized by repeated bottom-up and top-down processing. This design aimed to capture and consolidate features at multiple resolutions, enabling precise localization of keypoints across scales. Subsequent innovations, such as the High-Resolution Network (HRNet) [142] and its variants [18, 163], have further advanced spatial information preservation by maintaining high-resolution representations throughout the entire network. Recently, transformer-based methods [156, 165] have integrated attention mechanisms to capture long-range dependencies and spatial correlations among joints and appearances, expanding the possibilities of HPE. Beyond vision-based HPE, alternative sensors like inertial measurement units (IMUs) [48], millimeter-wave (mmWave) radar [126, 151, 149] and thermal cameras [130, 17] have been explored, tailoring HPE to diverse use cases. Additionally, multi-sensor and multi-modal datasets [86, 152, 3] and approaches [35, 172, 3] have emerged as an innovative step over unimodal HPE to overcome individual sensor and modality limitations, offering a synergistic fusion of data that promises increased robustness and precision, especially in scenarios marked by occlusions, changing lighting conditions, and dynamic environments.
1.2 Definition of Animal Pose Estimation
@@
APE, as discussed in this review, is defined as the task of identifying and localizing body joints of terrestrial mammals using a diverse array of sensor inputs. This task results in the reconstruction of the animal’s pose, depicted as either a skeletal representation marked by keypoints (Figure 1(a)) or as a comprehensive body mesh (Figure 1(b)).
(a) Keypoints Pose Representations. (a) A 2D keypoints pose representation example adapted from [100]. (b) A 3D keypoints pose representation example adapted from [53].
(b) Body Mesh Reconstructions. (a) Original RGB image. (b) Reconstructed mesh for (a) using the 3D model-based body mesh reconstruction adapted from [178]. (c) Target RGB image. (d) Reconstructed mesh for (c) applying the 3D model-free body mesh reconstruction method from MagicPony [148].
Figure 1: Different animal pose estimation methodologies.
@@
Figure 2: Transfering human pose estimation knowledge to proximal animal classes (e.g. Chimpanzees) for their pose estimation [124].
@@
Exploring APE through various sensory and modality inputs is important as animals may reside in diverse habitats which require different sensors for data collection. Most current APE methodologies are RGB-based, which predict the articulated joint positions from images or videos. However, RGB-based methods fail in many scenarios, such as nocturnal monitoring. For example, rodents are more active in the dark, and standard RGB cameras become inadequate due to poor lighting conditions. Consequently, infrared sensors have emerged as suitable tools for capturing the nocturnal activities of animals [145, 109]. Additionally, mmWave radar can detect movements but suffers from the scarcity of its generated point clouds; LiDAR adds value by providing depth information. Inertial sensors [108] can also track movement and orientation, offering invaluable insights but requiring on-animal attachment. In specific scenarios, even auditory cues [67] can indicate animal movement. This paper also explores the potential of using multiple sensors (i.e. RGB + acoustic) to improve APE performance.
@@
There are two representations for the final estimated animal poses regardless of our sensors or modalities of the APE input data. The first encompasses either two-dimensional (2D) or three-dimensional (3D) keypoint pose representations [55, 70, 100, 25] (Figure 1(a)). In this category, the animal’s anatomy is represented as a tree of interconnected keypoints. These keypoints present animal body joints, and their interconnections reflect the natural relationship between these joints, providing a skeletal representation of the animal’s posture. The second method, body mesh reconstruction [178, 121, 148, 153, 147], presents a 3D view of the animal’s morphology (Figure 1(b)). Within this framework, we distinguish between model-based and model-free reconstructions. The model-based technique (e.g. SMAL [178]) uses supervision from the predefined template shape of the creature learned from 3D toy animal scans. This template is then refined by optimizing the shape and pose parameters. In particular, the shape parameters capture variations in body metrics (e.g. proportions, height, and weight), while the pose parameters address observable deformations due to movement. The model-based approach can generate detailed models, but collecting the necessary data for new types of animals can be time-intensive, restricting their applicability to rare species. In contrast, the model-free approach operates without any supervision from the predefined template shape of the creature (e.g. MagicPony [148], DOVE [147] and 3D-Fauna [74]), which are more capable when estimating the animal poses for the diverse animal species.
@@
The transition from HPE to APE is based on HPE frameworks. Due to skeletal similarities between humans and certain animals like quadrupeds, HPE models for skeletal keypoint detection [142, 49, 21] and body mesh reconstruction [82] can be pertinent to APE, with modifications to cater to distinct anatomical and dynamic variations across animal species [70, 116, 178, 121]. One notable example is [124], which involves transferring human DensePose knowledge to proximal animal classes (Figure 2). This approach leverages the DensePose-COCO dataset, which contains dense human body annotations, and adapts it to animal species like chimpanzees. The transfer is facilitated by establishing correspondences between the 3D models of humans and animals using semantic alignment, so that the learned dense pose mappings for humans can be applied to the animal class. This enables pose extraction without dense animal-specific annotations. In addition, innovations in APE that allow adaptation to unseen animals and environmental changes can also improve the robustness of HPE, particularly in challenging scenarios like emergency response, where the surrounding environment changes rapidly. However, bridging these two fields requires adaptation of knowledge and consideration of the distinct challenges and ethical issues in human and animal studies. Ethical and responsible practices [59] are especially critical in HPE methods development, and similar approaches can be applied to the APE methods as well.
1.3 Challenges in Human and Animal Pose Estimation
@@
There are also unique challenges in APE. The animal kingdom has a much greater morphological diversity and inter/intra-species variations than humans. Different animal species exhibit significant domain discrepancies in physical characteristics, body structures, and postures. Even within the same species, there can be a large variety in body size, shape, texture, motion, and posture, presenting difficulties in creating universally applicable models [121]. For example, the pelage of porcupines can vary a lot, and this is often more pronounced and thicker in animals than in humans’ hair, which can make tracking skeleton segments tricky. Furthermore, many animals exhibit natural camouflage, travel in herds, interact frequently, or share similar appearances within their species [176]. These characteristics make it hard to distinguish individual animals and accurately estimate their poses as they blend into the background or appear similar to one another.
Environmental factors also play a pronounced role. Vegetation or terrain features in the natural environments may obstruct the view of animals. One example can be found in the AcinoSet dataset [53], where the cheetahs’ paws were excluded as the grass occluded them in most of the videos, making it difficult to estimate animal poses accurately. Moreover, changing lighting conditions can also heavily affect an animal’s appearance, especially when the sun’s position is moving in outdoor environments.
@@
Moreover, there are very limited annotated APE datasets [164]. Unlike HPE, which predominantly leans on supervised [142, 18] and semi-supervised learning methodologies [138, 110, 92] sustained by large-scale human pose datasets such as MSCOCO [75] and MPII [4], APE is hindered by a lack of such datasets, especially for endangered species and wildlife [176]. Also, it is often impractical to place animals in controlled environments to obtain ground truth poses, although zoos provide a semi-structured environment towards data collection. Therefore, unsupervised approaches that do not rely on labelled data have become popular in APE [148, 69].
1.4 Applications of Animal Pose Estimation
@@
Tracking: APE has been applied in various tracking contexts; for example, in the CATER [42] framework, APE is combined with environment reconstruction to track ants in complex natural environments, providing fine-scale motion trajectories even in visually challenging conditions. The SMART-BARN [98] system uses APE to track multiple animal species simultaneously in a large 3D arena and achieve accurate tracking and behavioural analysis in indoor and naturalistic environments. In the FaceMap [136] framework, APE is used for tracking the orofacial movements of mice, enabling detailed analysis of their facial expressions and related neural activities. For honey bees, APE tracked an entire colony over extended periods, capturing complex social behaviours and interactions without needing physical tags, enhancing our understanding of colony dynamics and individual behaviours [14].
Table 1: Table listing various applications of animal pose estimation. The paper citation is available in column Year.
@@
1 2019 Bain et al. Chimpanzee Tracking
2 2021 Bozek et al. Honey Bee Tracking
3 2022 I-MuPPET Pigeon Tracking
4 2023 CATER Desert Ants Tracking
5 2023 SMART-BARN Insects, Mammals, Tracking
Birds
6 2023 3D-POP Birds Tracking
7 2023 DeepWild Chimpanzees, Tracking
@@
8 2024 Facemap Mice Tracking
9 2024 3D-MuPPET Pigeon Tracking
10 2023 Stathopoulos et al. Sheep, Cows, Bear, Articulation
@@
11 2022 Sheppard et al. Mouse Behaviour
12 2023 Voloh et al. Macaque Behaviour
13 2023 BARN Macaque Behaviour
14 2023 Koger et al. Gelada Monkeys Behaviour
15 2023 AmadeusGPT Mammals Behaviour
16 2024 Testard et al. Macaque Behaviour
Articulation: The authors of [133] explore how APE can be used to understand animal articulation by predicting their 3D shapes from monocular images. They demonstrate that training a 2D keypoint estimator on a small set of labelled images (50-150) can generate pseudo-labels on a more extensive set of unlabeled web images. Then, these pseudo-labels are used to train models for 3D shape reconstruction of various articulated quadrupeds.
Behaviour Analysis: In agriculture, animal behaviour analysis based on APE plays an important role by enabling effective livestock management and health surveillance, which in turn provides continuous monitoring of movement and also early detection of animal’s illness, thereby ensuring optimal animal welfare [32, 122, 34, 37, 28, 33]. Recently, more APE methods have been applied to study the behaviour of more species, such as macaques [139, 155, 137], Gelada monkeys [61], and mice [127]. Since 2023, there has been a notable shift towards integrating language prompts with APE in behaviour analysis research, exemplified by developing models like AmadeusGPT [161]. This combination has opened up new possibilities for more comprehensive interpretations of animal behaviour.
1.5 Scope of this Survey